OpenStudy (nwolf):

f(x)-8x-7 find the slope and intercept

Directrix (directrix):

Is this f(x) = 8x - 7 OR perhaps f(x) = -8x - 7?

OpenStudy (nwolf):

I forgot the = so it is F(X)=-8x-7

OpenStudy (anonymous):

let f(x)=y f(x)=-8x-7 y=-8x-7 comparing this with y=mx+c here m=-8,c=-7 slope=-8,y-intercept(0,c)=(0,-7)

OpenStudy (nwolf):

so when they ask for the slope i put (0-7) and intercept?? 8

Directrix (directrix):

----------- The slope-intercept form of writing the equation of a line is the following: y = mx + b where m is the slope and b is the y-intercept. Compare your function to this form. y = -8x - 7 --> yours y = mx + b --> general form Comparing, you see that m = -8 and b = -7. Because b is the y-intercept, it has an x-coordinate of 0 which would give it the ordered pair (0, -7). Most people just say that the y-intercept is 7. But, if you need to use the y-intercept in a calculation, you would want to know how to write it as (0,-7).